OpenStudy (dumbcow):

I have a specific question. Can someone help clarify a solution for me (posted below) Thanks

OpenStudy (dumbcow):

where does the \[\left(\begin{matrix}100 \\ 50\end{matrix}\right)\] derive from?

OpenStudy (dumbcow):

sorry i should give the context, the question was find the avg value for sin^100 x i understand the process, but don't see how they obtain 100 choose 50 when the integral of every term is 0 for k not equal 0

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

the definition given of the integral of exp(ikx), says that if k is not equal to 0 then the integral is 0

OpenStudy (anonymous):

we must integrate sin(x)^100, that is [(exp(ix) - exp(-ix))/2i]^100 right?

OpenStudy (anonymous):

so... sin(x)^100 = [(exp(ix) - exp(-ix))/2i]^100 = (1/2^100)*(exp(ix) - exp(-ix))^100

OpenStudy (anonymous):

remember that i^100 = 1

OpenStudy (anonymous):

now, for this part: [(exp(ix) - exp(-ix))/2i]^100 use newton binomial formula

OpenStudy (anonymous):

[(exp(ix) - exp(-ix))/2i]^100 = \[\sum_{k=0}^{100} (-1)^{100-k} e^{ikx} e^{-i(100-k)x}\]

OpenStudy (anonymous):

grouping these exponentials in one, we get: \[\sum_{k=0}^{100} (-1)^{100-k} e^{i(2k+100)x}\]

OpenStudy (anonymous):

now check when 2k+100 is not equal to 0

OpenStudy (anonymous):

\[k \neq 50\]

OpenStudy (anonymous):

so... only when k = 50, the integral is equal to 2*pi, at that time, from the newton binomial formula is at \[\left(\begin{matrix}100 \\ 50\end{matrix}\right)\]

OpenStudy (anonymous):

so if we plug it all together, we find the result shown in the pic

OpenStudy (anonymous):

got it?

OpenStudy (dumbcow):

yep got it, didn't think to combine the terms thanks a lot